# Learning and Transferring Physical Models through Derivatives

### v2

@@ Abstract
We propose Derivative Learning (DERL), a supervised approach that models physical systems by learning their partial derivatives. We also leverage DERL to build physical models incrementally, by designing a distillation protocol that effectively transfers knowledge from a pre-trained model to a student one. We provide theoretical guarantees that DERL can learn the true physical system, being consistent with the underlying physical laws, even when using empirical derivatives. DERL outperforms state-of-the-art methods in generalizing an ODE to unseen initial conditions and a parametric PDE to unseen parameters. We also design a method based on DERL to transfer physical knowledge across models by extending them to new portions of the physical domain and a new range of PDE parameters. We believe this is the first attempt at building physical models incrementally in multiple stages.
@@ 1 Introduction
Machine Learning (ML) techniques have found great success in modeling dynamical and physical systems, including Partial Differential Equations (PDEs). The growing interest around this topic is driven by the many real-world problems that would benefit from accurate prediction of dynamical systems, such as weather prediction (Pathak et al., 2022), fluid modeling (Zhang et al., 2024), quantum mechanics (Mo et al., 2022), and molecular dynamics (Behler & Parrinello, 2007). These problems require grasping the essence of the system by modeling its evolution while following the underlying physical laws. Purely data-driven supervised models often fail at this task: while being able to approximate any function, they do not usually learn to maintain consistency with the physical dynamics of the system (Greydanus et al., 2019; Hansen et al., 2023), or fail to approximate it when only a few data points are available (Czarnecki et al., 2017). Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) emerged as an effective paradigm to learn the dynamics of a PDE, by explicitly including in the loss the PDE components evaluated using automatic differentiation (Baydin et al., 2018). This imposes physical consistency by design and allows the use of PINNs in regimes where data is scarce. Unfortunately, PINNs suffer from optimization issues (Wang et al., 2022a; b) which can lead to poor generalization (Wang et al., 2021).
@@
Our approach is based on distillation (Hinton et al., 2015), where the physical knowledge contained in a fixed trained model (the teacher) is transferred to a student model by distilling the teacher’s derivatives with DERL. We show that a distillation protocol based on DERL surpasses the performance of alternative protocols based on different physical information. We successfully transfer information to new portions of the domain, to new parameterization of the physical systems, and to longer time horizons (Section 5). In all cases, the resulting final model is able to solve the whole problem by only being trained on it in an incremental way. Figure 1 illustrates the aforementioned transfer protocol via distillation to new domains. To the best of our knowledge, we are the first to propose an approach to transfer knowledge across physical models to study their incremental design through distillation. More broadly, our transfer protocol can be applied in multiple stages, in a continual learning fashion (Parisi et al., 2019). In continual learning, a model is trained on a sequence of different tasks with the objective of learning new tasks without forgetting previous ones. Continual learning has never been studied on physical systems. Our experiments show that physical models exhibit some degree of forgetting, suggesting that applying existing continual learning strategies might help mitigate this issue on a long sequence of tasks. While we foresee the potential impact of continual learning for our case, in this paper, we focus on the transfer task through distillation (notably, a technique widely used also in continual learning (Li & Hoiem, 2016)) for 2 or 3 sequential steps. We leave a proper exploration of continual learning in physical systems to future work.
@@ 2 Background and Related Works
Classical numerical methods involve diverse strategies to solve ODEs and PDEs. While for the first, it might suffice to consider the integration in time ${\bm{u}}(t+\Delta t)={\bm{u}}(t)+\Delta t{\bm{f}}({\bm{u}}(t),t)$, PDEs often require more complex methodologies and assumptions to ensure the convergence to the true solution (Evans, 2022). A common strategy involves discretizing the variables on a grid to approximate derivatives, transforming the PDE into a system of ODEs (finite differences, Ferziger & Peric (2001)). Similarly, finite volumes methods (Ferziger & Peric, 2001) divide the space into small volumes, each one exchanging physical information and quantities with its neighbors via fluxes on their boundaries. Finally, the finite element method (Anderson et al., 2020) uses the weak formulation of PDEs, where the solution is tested against a base of functions on a mesh of points, transforming it into a complex system of ODEs. Numerical methods, on the other hand, often require accurately selecting the correct mesh, balancing refinement and complexity. In fact, they can also be very slow, as these methods need to invert quite large matrices. Hence, there has been a growing interest in ML models for PDEs as they are by far the fastest at inference and allow for easy queries of any point.
@@
Recurrent neural networks (Schmidt, 2019) are a prototypical example of systems that evolve through time based on their previous state, which makes them a good candidate to learn solutions to ODEs. However, they only work iteratively and are prone to error accumulation or the vanishing gradient effect (Pascanu et al., 2013). Similarly, ResNets (He et al., 2016) and NODEs (Chen et al., 2019) model the residual state update as a discrete or continuous derivative, but cannot predict complete trajectories at once or require an external ODE solver. Normalizing flows (Rezende & Mohamed, 2015) model dynamical systems such as particles by learning their distribution with iterative invertible mappings. Neural operators (Li et al., 2020b) and their evolution (Li et al., 2021) act as neural networks for entire functions by mapping initial conditions or parameters to a solution via kernel operators and Fourier transforms. While these methods are powerful and find many applications to real-world problems (Pathak et al., 2022), they require a large amount of data to generalize, are computationally intensive, and do not ensure that the solution is physically consistent. Neural Operators are also suited for a different purpose than ours, as they are not made to map a point in time-space $t,{\bm{x}}$ to its value ${\bm{u}}(t,{\bm{x}})$. Our work shares similarities with Sobolev learning (Czarnecki et al., 2017; Srinivas & Fleuret, 2018), which adds derivative learning terms to the supervised loss with application to reinforcement learning and machine vision. As Sobolev learning is not designed for physical systems, we adapted it by adding the IC and BC into its loss, including a term for both $\bm{u}$} and its derivatives $\mathrm{D}u$. We show that the pure derivative approach of DERL achieves a better performance.
@@
where $\mathcal{F}[\hat{u}$] is the evaluation of the PDE in equation 3 on the network $\hat{u}$. PINNs are used to increase the physical consistency of a model. However, PINNs suffer from optimization problems (Wang et al., 2021) and can fail to reach a minimum of the loss (Sun et al., 2020). To alleviate these issues, there exist methodologies to simplify the objective (Sun et al., 2020) or to choose the collocation points where the PDE residual is evaluated (Zhao, 2021; Lau et al., 2024). Both approaches are problem-specific and add complexity to the optimization process. Specific architectures for parametric PDEs are also developed (Zou & Karniadakis, 2023; Penwarden et al., 2023) but often require multiple instances to be trained. Sobolev norms have also been applied to PINNs for improving convergence guarantees (Son et al., 2021). Models inspired by physics exploit the formalisms of Hamiltonian (Greydanus et al., 2019) and Lagrangian (Cranmer et al., 2019) mechanics to be inherently consistent with the properties of the system, but they require the system to be conservative and to be (implicitly) formulated in the Lagrangian or Hamiltonian formalism, which is generally not the case. They also require an external solver to compute the trajectories.
@@
Distillation Hinton et al. (2015) was originally proposed as a method to transfer knowledge from a large teacher model to a small student model with minimal performance loss. Over time, distillation proved to be a useful tool for many cases Gu et al. (2023). Continual learning is one of them, where distillation is used to avoid forgetting by transferring the knowledge from a teacher trained on previous tasks to a student trained on the current one Li & Hoiem (2016); Carta et al. (2024). Existing reviews on knowledge transfer include (Gou et al., 2021; Yang et al., 2024) for an extensive treatment. To the best of our knowledge, we are the first to use distillation as an approach to transfer knowledge across physical models.
@@ 3.1 Theoretical analysis
The proof is in Appendix B.2. This result can be extended to include higher-order derivatives by adding the corresponding terms to the loss $L$. We use these theorems to prove that a neural network trained to optimize the loss in equation 6 learns the solution to a PDE. We prove the following theorem in Appendix B.3.
@@
Let $\bm{u}$} be a solution to a PDE of the form of equation 3, which can be time-dependent or time-independent. A Neural Network $\hat{{\bm{u}}}$ trained to optimize $L(\hat{{\bm{u}}},{\bm{u}})$ as in equation 6 converges to the solution of the PDE $\bm{u}$} and fulfills all three conditions.
When analytical derivatives are not available, we use empirical derivatives obtained via finite differences (Anderson et al., 2020): $\frac{\partial u}{\partial x_{i}}({\bm{x}})\simeq\frac{u({\bm{x}}+he_{i})-u({\bm{x}})}{h},$ where $e_{i}$ is the unit vector in the direction of $x_{i}$ and $h$ is a small positive real number. Appendix C shows that finite differences converge on the whole domain to the true derivatives when $h\rightarrow 0$. This guarantees that we can still approximate $u$ using empirical derivatives.
@@ 4 Learning Physics from Data
We validate DERL on a set of dynamical systems and physical PDEs, investigating the generalization capabilities on a novel set of domain points (section 4.1) and on a novel set of parameters for parametric equations (section 4.2). During the training phase, the models did not have access to the set of points and parameters used to evaluate their generalization. Table 1 summarizes the systems we consider, with additional results in appendix D. For each experiment, we compare $4$ approaches applied on a Multi-Layer Perceptron (MLP, Goodfellow et al. (2016)). Each of them learns the IC and the BC, together with a loss term for the domain $\Omega$ which is specific to the approach: (a) DERL (ours) optimizes equation 6, (b) Output Learning (OUTL) minimizes $\|u-\hat{u}\|_{L^{2}}$, (c) our extension of Sobolev learning (SOB) (Czarnecki et al., 2017) to dynamical systems include the loss term for both $u$ and $\mathrm{D}u$. For comparison, we also consider (d) OUTL+PINN, which consists of OUTL with an additional PINN loss term on the equation to improve its performance in the domain. The full description of the baselines and losses is in Appendix A. We evaluate all approaches with two metrics: (i) we measure the accuracy of the prediction by $\|u-\hat{u}\|_{2}$ - the $L^{2}$ distance between the true function $u$ and the estimate $\hat{u}$; (ii) considering MLPs as functions from the $\Omega$ to the domain of $u$, we measure how well a model satisfies the underlying physical equation $\mathcal{F}[u]=0$. To this end, we compute the $L^{2}$ norm of the PDE residual of the network $\|\mathcal{F}[\hat{u}]\|_{2}$, as typically done with PINNs.
@@ 4.1 Generalization to new domain points
The Allen-Cahn equation (AC in table 1) is a non-linear PDE in the domain $1,1]^{2}$. We consider $\lambda=0.01$ and the analytical solution $u=\sin(\pi x)\sin(\pi y)$. The BC, the external force $f$, and the partial derivatives are calculated from $u$ using the PDE. The MLPs, which predict $u(x,y)$ from $(x,y)$, have $4$ layers with $50$ units and $\tanh$ activation and are trained for $100$ epochs with the BFGS optimizer (Li & Fukushima, 2001) on $1000$ random collocation points. Additional details are provided in Appendix D.1.
Table 2 shows the relevant metrics for this task. DERL outperforms all the other approaches in learning the true solution $u$, surpassing OUTL by a factor of 10. Furthermore, DERL achieves a lower PDE residual than OUTL+PINN. This shows that learning the direct solution $u$ is not as effective as DERL, not even when leveraging PINNs as an additional source of information. DERL also outperforms SOB, showing that derivatives contain all the needed information, while adding other terms to the loss can interfere and deteriorate the performance. This is also an empirical confirmation of our theoretical results. We finally report that DERL is the approach that best approximates the true derivatives ($\|\mathrm{D}u-\mathrm{D}\hat{u}\|_{2}$), surpassing OUTL and OUTL+PINN by a factor of 5 and improving on SOB. This is, of course, expected as DERL’s loss target includes the derivatives. We provide additional results on this task in Appendix D.1.
@@
We consider the continuity equation (equation CO in table 1), a time-dependent PDE describing mass conservation of matter in a velocity field. The domain of the unknown density $u(t,x,y)$ is the 2D plane region $1.5,1.5]^{2}$ for $t\in[0,10]$. The velocity field $\bm{v}$} is given by ${\bm{v}}(x,y)=(-y,x)$. The density is null at the boundary and the IC is given by $4$ Gaussian densities (see Appendix D.2), as in Torres et al. (2024). The reference solution is calculated using the finite volumes method (Ferziger & Peric, 2001).The MLPs, which predict $u(t,x,y)$ from $(t,x,y)$, have the same architecture as the Allen-Cahn experiment and are trained for $200$ epochs with batch size $128$ and Adam optimizer (Kingma, 2014). Partial derivatives are empirical and were calculated by finite difference approximation. See D.2 for further details and results.
Table 2 reports the $L^{2}$ error on the solution $u$ and the PDE residual $L^{2}$ norm calculated on the finest grid available. DERL is the most effective method: the distance from the ground truth is second only to OUTL (although still comparable) and almost two times smaller than SOB. We conjecture this is due to its more complex optimization objective with possibly conflicting loss terms for $u$ and $\mathrm{D}u$. Plus, DERL is clearly best in terms of PDE residual, surpassing even OUTL+PINN, which means our model is more consistent with the underlying equation on domain points we have not trained the model on. On a side note, we report that the PINN loss alone fails to propagate the solution through time correctly, as reported in Appendix D.2. These results for PINNs are aligned with Wang et al. (2022a), where the authors conjectured that PINN’s gradients are biased towards high values of $t$ with the model failing to propagate information correctly from the IC.
@@
We now consider a well-known system of PDEs: 2D time-independent Navier-Stokes equations made of the momentum (NS.M) and incompressibility (NS.I) equations in Table 1. The domain is $(x,y)\in[-1,1]\times[-0.5,1.5]$ with the viscosity coefficient set to $\nu=\frac{1}{50}$. The particular solution we work with is the so-called Kovasznay flow (Drazin & Riley, 2009), representing the flow around a two-dimensional grid, with the analytical form given in appendix D.3. Training data is sampled from a regular grid of points with $dx=dy=0.01$, while testing data comes from a finer one, with $dx=dy=0.005$. We work with MLPs with $4$ layers of $64$ units and $\tanh$ activation, trained with the LBFGS optimizer for $100$ epochs. For this task, we also measure the error on vorticity $\mathbf{\omega}=\nabla\times{\bm{u}}=\sin x\sin ye^{-2\nu t}\hat{\mathbf{z}}$, which is a vector that describes the rotating motion of a fluid at a given point and is of particular interest for turbulent flows such as the one for the Navier-Stokes equation.
Table 3: Results for the Navier-Stokes Kovasznay flow experiment. We report the

### v3

@@ Abstract
We propose Derivative Learning (DERL), a supervised approach that models physical systems by learning their partial derivatives. We also leverage DERL to build physical models incrementally, by designing a distillation protocol that effectively transfers knowledge from a pre-trained model to a student one. We provide theoretical guarantees that DERL can learn the true physical system, being consistent with the underlying physical laws, even when using empirical derivatives. DERL outperforms state-of-the-art methods in generalizing an ODE to unseen initial conditions and a parametric PDE to unseen parameters. We also design a method based on DERL to transfer physical knowledge across models by extending them to new portions of the physical domain and a new range of PDE parameters. This introduces a new pipeline to build physical models incrementally in multiple stages.
@@ 1 Introduction
Machine Learning (ML) techniques have found great success in modeling dynamical and physical systems, including Partial Differential Equations (PDEs). The growing interest around this topic is driven by the many real-world problems that would benefit from accurate prediction of dynamical systems, such as weather prediction (Pathak et al., 2022), fluid modeling (Zhang et al., 2024), quantum mechanics (Mo et al., 2022), and molecular dynamics (Behler and Parrinello, 2007). These problems require grasping the essence of the system by modeling its evolution while following the underlying physical laws. Purely data-driven supervised models often fail at this task: while being able to approximate any function, they do not usually learn to maintain consistency with the physical dynamics of the system (Greydanus et al., 2019; Hansen et al., 2023), or fail to approximate it when only a few data points are available (Czarnecki et al., 2017). Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) emerged as an effective paradigm to learn the dynamics of a PDE, by explicitly including in the loss the PDE components evaluated using automatic differentiation (Baydin et al., 2018). This imposes physical consistency by design and allows the use of PINNs in regimes where data is scarce. Unfortunately, PINNs suffer from optimization issues (Wang et al., 2022a; b) which can lead to poor generalization (Wang et al., 2021).
@@
Our approach is based on distillation (Hinton et al., 2015), where the physical knowledge contained in a fixed trained model (the teacher) is transferred to a student model by distilling the teacher’s derivatives with DERL. We show that a distillation protocol based on DERL surpasses the performance of alternative protocols based on different physical information. We successfully transfer information to new portions of the domain, to new parameterization of the physical systems, and to longer time horizons (Section 5). In all cases, the resulting final model is able to solve the whole problem by only being trained on it in an incremental way. Figure 1 illustrates the aforementioned transfer protocol via distillation to new domains. To the best of our knowledge, this defines a new approach to transfer knowledge across physical models and to study their incremental design through distillation. More broadly, our transfer protocol can be applied in multiple stages, in a continual learning fashion (Parisi et al., 2019). In continual learning, a model is trained on a sequence of different tasks with the objective of learning new tasks without forgetting previous ones. Continual learning has never been studied on physical systems. Our experiments show that physical models exhibit some degree of forgetting, suggesting that applying existing continual learning strategies might help mitigate this issue on a long sequence of tasks. While we foresee the potential impact of continual learning for our case, in this paper, we focus on the transfer task through distillation (notably, a technique widely used also in continual learning (Li and Hoiem, 2016)) for 2 or 3 sequential steps.
@@ 2 Background and Related Works
Classical numerical methods involve diverse strategies to solve ODEs and PDEs. While for the first, it might suffice to consider the integration in time ${\bm{u}}(t+\Delta t)={\bm{u}}(t)+\Delta t{\bm{f}}({\bm{u}}(t),t)$, PDEs often require more complex methodologies and assumptions to ensure the convergence to the true solution (Evans, 2022). A common strategy involves discretizing the variables on a grid to approximate derivatives, transforming the PDE into a system of ODEs (finite differences, Ferziger and Peric (2001)). Similarly, finite volumes methods (Ferziger and Peric, 2001) divide the space into small volumes, each one exchanging physical information and quantities with its neighbors via fluxes on their boundaries. Finally, the finite element method (Anderson et al., 2020) uses the weak formulation of PDEs, where the solution is tested against a base of functions on a mesh of points, transforming it into a complex system of ODEs. To apply classical methods, one needs to specify a pre-defined grid or mesh of points where the solution is calculated, which cannot be changed afterwards. Although being precise and providing theoretical guarantees on the convergence to the true solution as the grid becomes finer and finer, numerical methods are limited by being tied to a specific mesh. To predict the values of the solution for other points, they need to recalculate the entire solution or to rely on approximations and interpolations. Hence, they require accurately selecting the correct mesh, balancing refinement and complexity. In fact, they can also be very slow, as these methods need to invert quite large matrices. Hence, there has been a growing interest in ML models for PDEs as they are by far the fastest at inference and allow for easy queries of any point. Their capabilities are limited only by the availability of data and, in general, fewer results on theoretical convergence.
@@
Recurrent neural networks (Schmidt, 2019) are a prototypical example of systems that evolve through time based on their previous state, which makes them a good candidate to learn solutions to ODEs. However, they only work iteratively and are prone to error accumulation or the vanishing gradient effect (Pascanu et al., 2013). Similarly, ResNets (He et al., 2016) and NODEs (Chen et al., 2019) model the residual state update as a discrete or continuous derivative, but cannot predict complete trajectories at once or require an external ODE solver. Normalizing flows (Rezende and Mohamed, 2015) model dynamical systems such as particles by learning their distribution with iterative invertible mappings. Neural operators (Li et al., 2020b) and their evolution (Li et al., 2021) act as neural networks for entire functions by mapping initial conditions or parameters to a solution via kernel operators and Fourier transforms. While these methods are powerful and find many applications to real-world problems (Pathak et al., 2022), they require a large amount of data to generalize, are computationally intensive, and do not ensure that the solution is physically consistent. Neural Operators are also suited for a different purpose than ours, as they are not made to map a point in time-space $t,{\bm{x}}$ to its value ${\bm{u}}(t,{\bm{x}})$. Our work shares similarities with Sobolev learning (Czarnecki et al., 2017; Srinivas and Fleuret, 2018), which adds derivative learning terms to the supervised loss with application to reinforcement learning and machine vision. As Sobolev learning is not designed for physical systems, we adapted it by adding the IC and BC into its loss, including a term for both $\bm{u}$} and its derivatives $\mathrm{D}u$. We show that the pure derivative approach of DERL achieves a better performance.
@@
where $\mathcal{F}[\hat{u}$] is the evaluation of the PDE in equation 3 on the network $\hat{u}$. PINNs are used to increase the physical consistency of a model. However, PINNs suffer from optimization problems (Wang et al., 2021) and can fail to reach a minimum of the loss (Sun et al., 2020). To alleviate these issues, there exist methodologies to simplify the objective (Sun et al., 2020) or to choose the collocation points where the PDE residual is evaluated (Zhao, 2021; Lau et al., 2024). Both approaches are problem-specific and add complexity to the optimization process. Specific architectures for parametric PDEs are also developed (Zou and Karniadakis, 2023; Penwarden et al., 2023) but often require multiple instances to be trained. Sobolev norms have also been applied to PINNs for improving convergence guarantees (Son et al., 2021). Models inspired by physics exploit the formalisms of Hamiltonian (Greydanus et al., 2019) and Lagrangian (Cranmer et al., 2019) mechanics to be inherently consistent with the properties of the system, but they require the system to be conservative and to be (implicitly) formulated in the Lagrangian or Hamiltonian formalism, which is generally not the case. They also require an external solver to compute the trajectories.
@@
Distillation Hinton et al. (2015) was originally proposed as a method to transfer knowledge from a large teacher model to a small student model with minimal performance loss. Over time, distillation proved to be a useful tool for many cases Gu et al. (2023). Continual learning is one of them, where distillation is used to avoid forgetting by transferring the knowledge from a teacher trained on previous tasks to a student trained on the current one Li and Hoiem (2016); Carta et al. (2024). Existing reviews on knowledge transfer include (Gou et al., 2021; Yang et al., 2024) for an extensive treatment. We employ distillation as a means of enabling knowledge transfer across physical models.
@@ 3.1 Theoretical analysis
The proof is in Appendix B.2. This result can be extended to include higher-order derivatives by adding the corresponding terms to the loss $L$, leading to a convergence in $W^{m,2}$, where $m$ is the highest order of derivative included. We use these theorems to prove that a neural network trained to optimize the loss in equation 6 learns the solution to a PDE. We prove the following theorem in Appendix B.3.
@@
Let $\bm{u}$} be a solution to a PDE of the form of equation 3 with $\mathcal{F}$ continuous, and let $m$ be the order of the PDE, that is, the maximum order of derivative of $u$ that appears in $\mathcal{F}[{\bm{u}}]$. A Neural Network $\hat{{\bm{u}}}$ trained with the loss in equation 6, with derivatives terms up to order $m$, converges to the solution of the PDE $\bm{u}$} and fulfills all three conditions.
Remark 3.4.
PINNs incorporate the information on the PDE directly through the unsupervised PDE residual term $\|\mathcal{F}[\hat{{\bm{u}}}]\|_{2}^{2}$ in the loss. This way, PINNs have direct access to the definition of the physical system and the underlying equation representing its dynamics. On the contrary, DERL only learns the derivatives of the true solution $\bm{u}$} through the derivative learning term in equation 6, without seeing the true PDE. This theorem shows that learning the derivatives is enough for the DERL to recover a solution $\hat{{\bm{u}}}$ that fulfills the PDE, that is $\mathcal{F}[\hat{{\bm{u}}}]\rightarrow 0$. Hence, DERL learns to be consistent with the physical system without seeing the underlying equation. To provide an explicit example, consider a simple PDE such as $\mathcal{F}[u]=u_{t}+u_{x}=0$ and let $u$ be a solution related to some specified initial and boundary conditions. As the loss in equation 6 converges to zero, we have that $\hat{u}_{t}\rightarrow u_{t}$ and $\hat{u}_{x}\rightarrow u_{x}$ as per Theorem 3.2. Hence, $\mathcal{F}[\hat{u}]=\hat{u}_{t}+\hat{u}_{x}\rightarrow u_{t}+u_{x}=\mathcal{F}[u]=0$.
Remark 3.5.
While perfect consistency requires enough derivative terms up to the order of the PDE, our results show that, in practice, the loss in equation 6 is enough for the majority of cases, already providing better results than the other baselines. This clearly shows that learning just $\bm{u}$}, as in most supervised approaches, while derivatives already provide sufficient and richer information on the system.
When analytical derivatives are not available, we use empirical derivatives obtained via finite differences (Anderson et al., 2020): $\frac{\partial u}{\partial x_{i}}({\bm{x}})\simeq\frac{u({\bm{x}}+he_{i})-u({\bm{x}})}{h},$ where $e_{i}$ is the unit vector in the direction of $x_{i}$ and $h$ is a small positive real number. These approximations introduce errors, but the following result ensures that with a small enough $h$, these derivatives are very similar to the true ones. The proof is in Appendix B.4.
Theorem 3.6.
where ${\bm{e}}_{i}$ is the unit vector in the $x_{i}$ direction. Then, we have that $\|\mathrm{D}^{\epsilon}_{x_{i}}u\|_{L^{2}(\mathbb{R}^{n})}\leq\|\frac{\partial u}{\partial x_{i}}\|_{L^{2}(\mathbb{R}^{n})}$ and $\mathrm{D}^{\epsilon}_{x_{i}}u\rightarrow u_{x_{i}}$ in $L^{2}(\mathbb{R}^{n})$, which means that empirical derivatives converge to weak (or true) ones as $\epsilon\rightarrow 0$. The same results hold for any open set $\Omega$, with the convergence being true on every compact subset of $\Omega$. In practical terms, empirical derivatives converge a.e. for every point distant at least $h$ from the boundary.
Remark 3.7.
The approximation used to calculate the derivative is of first order $O(h)$. If sufficient data is available, one can further improve precision by considering higher-order approximations, such as the central scheme (Anderson et al., 2020), which has an $O(h^{2})$ error, or beyond.
@@ 4 Learning Physics from Data
We validate DERL on a set of dynamical systems and physical PDEs [^1] The code to reproduce the experiments is available at https://github.com/AlexThirty/DERL, investigating the generalization capabilities on a novel set of domain points (section 4.1) and on a novel set of parameters for parametric equations (section 4.2). During the training phase, the models did not have access to the set of points and parameters used to evaluate their generalization. Table 1 summarizes the systems we consider, with additional results in appendix C. For each experiment, we compare $4$ approaches applied on a Multi-Layer Perceptron (MLP, Goodfellow et al. (2016)). Each of them learns the IC and the BC, together with a loss term for the domain $\Omega$ which is specific to the approach: (a) DERL (ours) optimizes equation 6, (b) Output Learning (OUTL) minimizes $\|u-\hat{u}\|_{L^{2}}$, (c) our extension of Sobolev learning (SOB) (Czarnecki et al., 2017) to dynamical systems include the loss term for both $u$ and $\mathrm{D}u$. For comparison, we also consider (d) OUTL+PINN, which consists of OUTL with an additional PINN loss term on the equation to improve its performance in the domain. The full description of the baselines and losses is in Appendix A. We evaluate all approaches with two metrics: (i) we measure the accuracy of the prediction by $\|u-\hat{u}\|_{2}$ - the $L^{2}$ distance between the true function $u$ and the estimate $\hat{u}$; (ii) considering MLPs as functions from the $\Omega$ to the domain of $u$, we measure how well a model satisfies the underlying physical equation $\mathcal{F}[u]=0$. To this end, we compute the $L^{2}$ norm of the PDE residual of the network $\|\mathcal{F}[\hat{u}]\|_{2}$, as typically done with PINNs.
@@ 4.1 Generalization to new domain points
The Allen-Cahn equation (AC in table 1) is a non-linear PDE in the domain $1,1]^{2}$. We consider $\lambda=0.01$ and the analytical solution $u=\sin(\pi x)\sin(\pi y)$. The BC, the external force $f$, and the partial derivatives are calculated from $u$ using the PDE. The MLPs, which predict $u(x,y)$ from $(x,y)$, have $4$ layers with $50$ units (similar to Raissi et al. (2019)) and $\tanh$ activation and are trained for $100$ epochs with the BFGS optimizer (Li and Fukushima, 2001) on $1000$ random collocation points. Additional details are provided in Appendix C.1.
Table 2 shows the relevant metrics for this task. DERL outperforms all the other approaches in learning the true solution $u$, surpassing OUTL by a factor of 10. Furthermore, DERL achieves a lower PDE residual than OUTL+PINN. This shows that learning the direct solution $u$ is not as effective as DERL, not even when leveraging PINNs as an additional source of information. DERL also outperforms SOB, showing that derivatives contain all the needed information, while adding other terms to the loss can interfere and deteriorate the performance. This is also an empirical confirmation of our theoretical results. We finally report that DERL is the approach that best approximates the true derivatives ($\|\mathrm{D}u-\mathrm{D}\hat{u}\|_{2}$), surpassing OUTL and OUTL+PINN by a factor of 5 and improving on SOB. This is, of course, expected as DERL’s loss target includes the derivatives. We provide additional results on this task in Appendix C.1.
@@
We consider the continuity equation (equation CO in table 1), a time-dependent PDE describing mass conservation of matter in a velocity field. The domain of the unknown density $u(t,x,y)$ is the 2D plane region $1.5,1.5]^{2}$ for $t\in[0,10]$. The velocity field $\bm{v}$} is given by ${\bm{v}}(x,y)=(-y,x)$. The density is null at the boundary and the IC is given by $4$ Gaussian densities (see Appendix C.2), as in Torres et al. (2024). The reference solution is calculated using the finite volumes method (Ferziger and Peric, 2001).The MLPs, which predict $u(t,x,y)$ from $(t,x,y)$, have the same architecture as the Allen-Cahn experiment and are trained for $200$ epochs with batch size $128$ and Adam optimizer (Kingma, 2014). Partial derivatives are empirical and were calculated by finite difference approximation. See C.2 for further details and results.
Table 2 reports the $L^{2}$ error on the solution $u$ and the PDE residual $L^{2}$ norm calculated on the finest grid available. DERL is the most effective method: the distance from the ground truth is second only to OUTL (although still comparable) and almost two times smaller than SOB. We conjecture this is due to its more complex optimization objective with possibly conflicting loss terms for $u$ and $\mathrm{D}u$. Plus, DERL is clearly best in terms of PDE residual, surpassing even OUTL+PINN, which means our model is more consistent with the underlying equation on domain points we have not trained the model on. On a side note, we report that the PINN loss alone fails to propagate the solution through time correctly, as reported in Appendix C.2. These results for PINNs are aligned with Wang et al. (2022a), where the authors conjectured that PINN’s gradients are biased towards high values of $t$ with the model failing to propagate information correctly from the IC.
@@
We now consider a well-known system of PDEs: 2D time-independent Navier-Stokes equations made of the momentum (NS.M) and incompressibility (NS.I) equations in Table 1. The domain is $(x,y)\in[-1,1]\times[-0.5,1.5]$ with the viscosity coefficient set to $\nu=\frac{1}{50}$. The particular solution we work with is the so-called Kovasznay flow (Drazin and Riley, 2009), representing the flow around a two-dimensional grid, with the analytical form given in appendix C.3. Training data is sampled from a regular grid of points with $dx=dy=0.01$, while testing data comes from a finer one, with $dx=dy=0.005$. We work with MLPs with $4$ layers of $64$ units and $\tanh$ activation, trained with the LBFGS optimizer for $100$ epochs. Here, we also measure the error on vorticity $\mathbf{\omega}=\nabla\times{\bm{u}}=\sin x\sin ye^{-2\nu t}\hat{\mathbf{z}}$, which is a vector that describes the rotating motion of a fluid at a given point and is of particular interest for turbulent flows such as this experiment.
Table 3: Results for the Navier-Stokes Kovasznay flow experiment.